OpenStudy (anonymous):

Landon opened a savings account with a deposit of $3,786.45. Nine years later, the balance in the account is $5,135.29. If the interest rate is 3.4%, how often does the interest compound?

OpenStudy (anonymous):

\[A=P\left(1+{r\over100}\right)^n\] A=5135.29 P=3768.45 r=3.4 n=? plug in and solve.

OpenStudy (anonymous):

ooh.. they want "how often"... k

OpenStudy (anonymous):

my bad.. n = 9 and we use \[A=P\left(1+{r\over 100k}\right)^{kn}\]

OpenStudy (anonymous):

what would i use for k ?

OpenStudy (anonymous):

\[5135.29=3768.45\left(1+{3.4\over100k}\right)^{9k}\\ {5135.29\over3768.45}=\left(1+{0.034\over k}\right)^{9k}\\ 1.3627=\left(1+{0.034\over k}\right)^{9k}\\ \log_{10}(1.3627)=9k\times\log_{10}\left(1+{0.034\over k}\right)\\ 0.1344=9k\times\log_{10}\left(1+{0.034\over k}\right) \]

OpenStudy (anonymous):

the only way I see is by trial and error 1) usually savings accounts are coumponded monthly.. so, check for k = 12 2) if does not work, check semiannually k=2 3) check annually. k=1

OpenStudy (anonymous):

alright thanks

OpenStudy (anonymous):

the answer must be "ANNUALLY" , with k=1